A boy has a mass of 50 kg. He jumps at a speed of 2.5 m/s into a stationary boat which has a mass of 70 kg. Please calculate their speed (ignore friction).


1
Expert's answer
2022-01-31T07:33:43-0500

Given:

"m_1=50\:\rm kg"

"v_1=2.5\:\rm m/s"

"m_2=70\:\rm kg"

"v_2=0\:\rm m/s"


The law of conservation of momentum says

"m_1v_1+m_2v_2=(m_1+m_2)u"

The final speed of a boat with a boy

"u=\frac{m_1v_1+m_2v_2}{m_1+m_2}=\frac{50*2.5+70*0}{50+70}=1.0\:\rm m/s"